# Student-AI Interaction: A Case Study of CS1 students

### v1

@@ 4.7. Participants’ Self-efficacy
To evaluate the potential impact of Generative AI on self-efficacy, we conducted pre-study and post-study self-efficacy surveys. Unfortunately, for the first five participants (P1 to P5), we did not capture the pre-study self-efficacy data. Figure 3 presents the participants’ perception of self-efficacy before and after using Generative AI. Figure 3(a) shows the distribution of perception of self-efficacy before and after the study. Given the small number of observations, we cannot draw any statistically-sound conclusion about any difference between the distributions.
@@
In contrast, self-efficacy in P9 decreased from 2.58 before the study to 1.75 after the study. Although P9 completed all programming tasks successfully, she copied the full description to the plug-in, asked ChatGPT for the answer, and pasted the results as the submission. Similarly, this occurred for P13, who began with a pre-study self-efficacy score of 3.92, however, the post-study mean score dropped to 3.00. He submitted an incorrect solution for one task and did not finish the other two tasks, she used Generative AI in a step-by-step approach.
@@ 5. Threats to Validity
Replicability Can others replicate our results? Because Generative AI and its usage and acceptance are rapidly emerging, we do not know if the results of this study will be replicable due to changing student behaviors and rapid advances in Generative AI. We encourage other researchers to replicate our study for different populations to check for emerging behaviors of students in using Generative AI.
@@ 6.1. Interactions with Generative AI
In this study, our objective was to investigate how CS1 students utilize ChatGPT for programming questions to assess whether Generative AI serves as a help-seeking tool for beginners in programming. As observed in other recent studies  (Kazemitabaar et al., 2023b; Prather et al., 2024), our findings indicate that a large number of Generative AI users rely on providing full descriptions of programming questions to find solutions without making sufficient effort on their own, even under supervision. This trend aligns with the patterns of over-reliance on LLMs identified by Kazemitabaar et al.  (Kazemitabaar et al., 2023b), particularly among novices using the ’AI single prompt’ approach, which resulted in lower performance on subsequent tasks. The observed behavior raises concerns about the potential overreliance on Generative AI in educational settings, where students might increasingly rely on Generative AI to provide all solutions to the detriment of their learning. This echoes the observations of Fernandez and Cornell  (Fernandez and Cornell, 2024), who emphasized the need for careful integration of AI-driven code generation tools to avoid such overreliance. Help-seeking is crucial for students to grasp new concepts, acquire skills, and tackle challenges in their computing courses (Hou et al., 2024). However, when participants use full descriptions of programming tasks as prompts and accept complete Generative AI-generated answers, Generative AI may not effectively fulfill its role as a help-seeking tool that constructively aids struggling students, but an oracle that does the learners’ job for them. This concern was also raised by Jošt et al. (Jošt et al., 2024), highlighting the need for instructional strategies that emphasize breaking down problems and leveraging Generative AI for incremental learning. Our findings, alongside those of Prather et al.  (Prather et al., 2024), suggest that some students may struggle with new metacognitive difficulties when using generative AI tools, including being conceptually behind in course material but unaware of it due to a false sense of confidence. Further exploration is needed to understand the underlying reasons for this undesired behavior and to encourage a more constructive use of Generative AI that promotes deep understanding and problem-solving skills in computing education.
Our observations reveal two main types of behavior: iterative and linear. Iterative users refined their prompts to achieve correct answers, while linear users used full descriptions to find answers directly. This behavioral split reflects the findings of Vadaparty et al. (Vadaparty et al., 2024), who noted similar patterns in student interactions with Generative AI in a CS1 course. The iterative approach can improve learning by encouraging deeper engagement with problem-solving processes, whereas the linear approach may indicate a tendency to seek quick fixes.
@@
Our results show that in 30% of prompts, users accepted the Generative AI response, while in 70%, AI responses were used to learn concepts. This dual role of Generative AI as both an answer provider and a personal tutor aligns with the observations by Prather et al. (Prather et al., 2024) regarding the mixed impact of Generative AI on student learning, where Generative AI facilitated both understanding and dependency. Effective Generative AI integration should balance assistance with promoting independent problem-solving skills.
We can categorize the nature of interactions with Generative AI based on the reliance of the prompt on the prior prompts. into two groups (1) stateless and (2) stateful. In the stateless interactions, individual prompts are independent of the prior prompts, hence prompts can be interpreted and answered independently. However, in stateful interactions, the prompt assumes that the Generative AI uses the history of the user’s interaction. For instance, P7 used stateful interaction, where he asked “So, what was wrong with my initial code?”.
We observed that seven participants P4, P7, P9, P10, P13, P14, and P15 employed the same approach on all tasks. We call their strategy for problem-solving, single-approach strategy. In contrast, P1, P2, P3, P5, P6, P8, P11, and P12 used different approaches for different tasks that we call mixed-approach.
As generative AI becomes more commonplace and more learners can access them for diverse sets of topics, a more granular investigation of these approaches becomes more important to guide pedagogy.
6.2. Self Efficacy and Generative AI
Our observations suggest that users who used full descriptions of questions may prefer smooth interactions without challenges. This behavior correlates with lower self-efficacy scores, similar to the findings of Xue et al. (Xue et al., 2024), who observed that students with lower self-efficacy tend to rely more heavily on Generative AI tools. Overall, most users’ self-efficacy levels increased after programming with Generative AI. This suggests the varied impacts of integrating Generative AI in educational contexts, shaping self-efficacy outcomes according to individual learning strategies and initial confidence levels. Enhancing self-efficacy through scaffolded Generative AI interactions could support more confident and independent problem-solving.
@@ 6.3. Implications for Computing Education
The current approach to prompt engineering focuses on productivity, aiming to help developers find final solutions quickly. However, in an educational context, the goal is not merely to reach a solution, but to ensure that students achieve the learning objectives. This approach is inherently different from current practices in prompt engineering, as proper a learning strategy may require expecting or even encouraging a significant amount of purposeful failures along the way, while productivity aims to minimize and eradicate those failures. Thus, it is necessary to investigate prompt engineering approaches that prioritize individual learning.
Our results suggest that, in the physical presence of a researcher and with knowledge of being recorded, in a considerable number of cases, the participants directly resorted to Generative AI to provide solutions to the tasks without making any attempts on their own. This undesirable way of help-seeking, if used as the default approach in solving homework problems, can negatively impact students’ learning. This suggests the need for students to improve their self-regulation skills, e.g. self-monitoring (Schunk and Zimmerman, 2012) to monitor and reflect the intensity and frequency of their use of Generative AI. Similarly, Generative AI educational tool builders should consider supporting such strategies that enhance students’ self-regulation and help-seeking behavior when they use Generative AI.
One can imagine that a possible remedy for this problem is to devise novel homework assignments that can be difficult for Generative AI to solve. However, as Generative AI tools become more sophisticated and powerful, the arms race between educators and Generative AI seems to be a losing proposition, especially in introductory courses such as CS1 that only include basic algorithmic thinking that there is an upper limit to the appropriate complexity of problems. To discourage students from using prompts to find complete solutions to homework problems, we should investigate novel types of problems that are compatible with the era of Generative AI. Recent work such as Promptly (Denny et al., 2024a) that uses the graphical representation for problem description instead of a textual description is a step in this direction.

### v2

@@ 4.3. RQ3: How does student self-efficacy change before and after programming with Generative AI?
Figure 3 presents the participants’ perception of self-efficacy before and after using Generative AI. Figure 3(a) shows the distribution of perception of self-efficacy before and after the study. Given the small number of observations, we cannot draw any statistically-sound conclusion about any difference between the distributions.
@@
In contrast, self-efficacy in P9 decreased from $2.58$, before the study, to $1.75$ after the study. Although P9 completed all programming tasks successfully, she copied the full description to the plug-in, asked ChatGPT for the answer, and pasted the results as the submission. Similarly, this occurred for P13, who began with a pre-study self-efficacy score of $3.92$, however, the post-study mean score dropped to 3.00. He submitted an incorrect solution for one task and did not finish the other two tasks, she used Generative AI in a step-by-step approach.
@@ 5. Threats to Validity
Replicability Can others replicate our results? Because Generative AI and its usage and acceptance are rapidly emerging, we do not know if the results of this study will be replicable due to changing student behaviors and rapid advances in Generative AI. We encourage other researchers to replicate our study for different populations to check for emerging behaviors of students when using Generative AI.
@@ 6.1. Interactions with Generative AI
In this study, our objective was to investigate how CS1 students utilize Generative AI for programming questions to assess whether Generative AI serves as a help-seeking tool for beginners in programming. As observed in other recent studies  (Kazemitabaar et al., 2023c; Prather et al., 2024b), our findings indicate that a large number of Generative AI users rely on providing full descriptions of programming questions to find solutions without making sufficient effort on their own, even under supervision. This trend aligns with the patterns of over-reliance on LLMs identified by Kazemitabaar et al. (Kazemitabaar et al., 2023c), particularly among novices using the ’AI single prompt’ approach, which resulted in lower performance on subsequent tasks. The observed behavior raises concerns about the potential overreliance on Generative AI in educational settings, where students might increasingly rely on Generative AI to provide all solutions to the detriment of their learning. This echoes the observations of Fernandez and Cornell  (Fernandez and Cornell, 2024), who emphasized the need for careful integration of AI-driven code generation tools to avoid such overreliance. Help-seeking is crucial for students to grasp new concepts, acquire skills, and tackle challenges in their computing courses (Hou et al., 2024b). However, when participants use full descriptions of programming tasks as prompts and accept complete Generative AI-generated responses, Generative AI may not effectively fulfill its role as a help-seeking tool that constructively aids struggling students, but an oracle that does the learners’ job for them. This concern was also raised by Jošt et al.(Jošt et al., 2024), highlighting the need for instructional strategies that emphasize breaking down problems and leveraging Generative AI for incremental learning. Our findings, alongside those of Prather et al. (Prather et al., 2024b), suggest that some students may struggle with new metacognitive difficulties when using generative AI tools, including being conceptually behind in course material but unaware of it due to a false sense of confidence. Further exploration is needed to understand the underlying reasons for this undesired behavior and to encourage a more constructive use of Generative AI that promotes deep understanding and problem-solving skills in computing education.
Our observations reveal two main types of behavior in problem solving: iterative and linear. Students who employed an iterative approach refined their prompts to achieve correct answers, while students who used a linear approach used the full description of the problems to find answers directly. This behavioral split reflects the findings of Vadaparty et al. (Vadaparty et al., 2024), who noted similar patterns in student interactions with Generative AI in a CS1 course. The iterative approach can improve learning by encouraging deeper engagement with problem-solving processes, whereas the linear approach may indicate a tendency to seek quick fixes.
@@
Our results show that in 30% of prompts, users accepted the Generative AI response, while in 70%, AI responses were used to learn concepts. This dual role of Generative AI as both an answer provider and a personal tutor aligns with the observations by Prather et al. (Prather et al., 2024b) regarding the mixed impact of Generative AI on student learning, where Generative AI facilitated both understanding and dependency. Effective Generative AI integration should balance assistance with promoting independent problem-solving skills.
We can categorize the nature of interactions with Generative AI based on the reliance of a prompt on prior prompts into two groups: (1) stateless and (2) stateful. In the stateless interactions, individual prompts are independent of the prior prompts, hence prompts can be interpreted and answered independently. However, in stateful interactions, the prompt assumes that Generative AI uses the history of the user’s interaction. For instance, P7 used stateful interaction, where he asked “So, what was wrong with my initial code?”.
We observed that seven participants P4, P7, P9, P10, P13, P14, and P15 employed the same approach on all tasks. We call their strategy for problem-solving, single-approach strategy. In contrast, P1, P2, P3, P5, P6, P8, P11, and P12 used different approaches for different tasks that we call mixed-approach. As generative AI becomes more commonplace and more learners can access them for diverse sets of topics, a more granular investigation of these approaches becomes more important to guide pedagogy.
6.2. Self-efficacy and Generative AI
Our observations suggest that users who used full descriptions of the questions may prefer smooth interactions without challenges. This behavior correlates with lower self-efficacy scores, similar to the findings of Xue et al. (Xue et al., 2024) and Prather et al. (Prather et al., 2024b), who observed that students with lower self-efficacy tend to rely more heavily on Generative AI tools. Overall, some users’ self-efficacy levels increased after programming with Generative AI. This suggests the varied impacts of integrating Generative AI in educational contexts, shaping self-efficacy outcomes according to individual learning strategies and initial confidence levels. Enhancing self-efficacy through scaffolded Generative AI interactions could support more confident and independent problem-solving.
@@ 6.3. Implications for Computing Education
The current advancement in prompt engineering focuses on productivity, aiming to help developers find final solutions quickly. However, in an educational context, the goal is not merely to reach a solution, but to ensure that students achieve the learning objectives. This goal is inherently different from prompt engineering for productivity improvement, which aims to minimize and eradicate failures, a proper learning strategy may require expecting or even encouraging a significant amount of purposeful failures along the way. Thus, prompt engineering research prioritizing individual learning, as seen in works like Jin et al. (Jin et al., 2024), should be further investigated.
Our results suggest that, even in the physical presence of a researcher and with knowledge of being recorded, in a considerable number of cases, the participants directly resorted to Generative AI to provide solutions to the tasks without making any attempts on their own. This undesirable way of help-seeking, if used as the default approach in solving homework problems, can negatively impact students’ learning. This suggests the need for students to improve their self-regulation skills, e.g. self-monitoring (Schunk and Zimmerman, 2012) to monitor and reflect the intensity and frequency of their use of Generative AI. Similarly, Generative AI educational tool builders should consider supporting such strategies that enhance students’ self-regulation and help-seeking behavior when they use Generative AI.
The findings of Margulieux et al. (Margulieux et al., 2024) are particularly relevant here. In their study, they found that some students used Generative AI to support and not replace their critical thinking and problem solving. However, they also noticed that the weakest students tended to use Generative AI earlier in the problem solving process. Their findings mirror those in our study. However, Margulieux et al.were somewhat optimistic that their findings meant that users who needed help and support could receive it. Our findings, as well as those by Prather et al. (Prather et al., 2024b), show that lower-performing students using Generative AI earlier in the problem solving process is likely an indication of overreliance and a failure to properly self-regulate with these tools.
One can imagine that a possible remedy for this problem is to devise novel homework assignments that can be difficult for Generative AI to solve. However, as Generative AI tools become more sophisticated and powerful, the arms race between educators and Generative AI seems to be a losing proposition, especially in introductory courses such as CS1 that only include basic algorithmic thinking that there is an upper limit to the appropriate complexity of problems  (Prather et al., 2023a; Denny et al., 2024c). To discourage students from using prompts to find complete solutions to homework problems, we should investigate novel types of problems that are compatible with the era of Generative AI. Recent work such as Prompt Problems (Denny et al., 2024a) that uses the graphical representation for problem description instead of a textual description is a step in this direction. Even though the newest models with visual modality have been able to solve Prompt Problems (Hou et al., 2024a), they remain useful as a way to scaffold students usage of Generative AI by helping them learn problem decomposition, iterative problem solving, describing a problem, and prompt engineering (Prather et al., 2024a).